The amounts of pollutants A entering the atmosphere is directly proportional to the population P living in an area. If 25000 tons of pollutant enter the atmosphere in an with a population of 17750, how many tons of pollutants enter the atmosphere in manila with a population of 2450500?
Answer by mananth(16946)

The amount of pollutants A entering the atmosphere is directly proportional to the population P living in an?
area .If 25 000 tons of pollutants enter the atmosphere in an area with a population of 17 750, how many tons of pollutants enter the atmosphere in Manila with a population of 2 450 500?
with solutions please
...
A alpha P
A=KP where K is a constant.
...
25 000 tons of pollutants enter the atmosphere in an area with a population of 17 750,
25,000=K*17,750
K= 25,000/17,750
K=1.408
...
Manila with a population of 2 450 500
P=2 450 500
K= 1.408
...
A= 1.408*2,450,500
A=3,450,304 Tons
...
m.ananth@hotmail.ca